# Title keyword 5 for citation Title for citation Title for citation

### v2

@@ 3.2.4 AI-Based Mobility Management
Mobility management is a critical component in modern wireless communication networks. Mobility management includes mobility in the idle state and mobility in the connected state. Due to the increasing number of mobile users and the increasing complexity of services, how to effectively manage mobility to improve network performance and user experience has become a critical issue.
Traditional mobility management relies heavily on preset policies and rules. In the idle state, mobility management involves cell selection and switching. Cell selection is usually based on signal strength, network load, and priority. In the connected state, the device must decide whether to switch cells based on network conditions, such as network coverage, quality, and load. In addition, the device may also undergo same-frequency switching or cross-frequency switching. Frequency switching refers to the equipment at the same frequency switching between two villages; different frequency switching equipment is involved in switching between the different frequencies of the villages.
However, there are some bottlenecks in the traditional mobility management approach. One is that these techniques might not manage user interactions fast enough when the number of mobile users spikes, leading to excessive network traffic and poor performance. The second is that typical mobility management techniques may be challenging in complex network situations, such as big public spaces or events, which could cause network congestion. Furthermore, users traveling at fast speeds—like those in cars—may have unreliable network connections due to frequent switching.
Over the past few decades, many schemes have been proposed to optimize mobility management [90], [91], [92], [93]. They all have a reactive nature; adjustments are often made after rather than pre-planning when dealing with mobility issues. This design can lead to delays and inefficiencies in mobility management and some unnecessary waste of resources.
To address these problems, AI has been introduced into mobility management. The advantage of AI is that it can process large amounts of data, learn automatically from data, and optimize strategies without human intervention. This allows AI to manage mobility more effectively, which improves network performance and user experience. Starting as early as 2015, 3GPP began to study how AI technology could be integrated into 5G networks. For example, the 3GPP SA2 and SA5 working groups’ ”Enablers of Network Automation” and ”Enhancement of Managed Data Analytics Services” projects made impressive progress of [94].
AI has shown significant advantages in mobility management with its superior fitting capabilities. On the one hand, using techniques such as deep learning, AI can predict user behavior, such as movement speed and direction, so that cell selection and switching can be done in advance to reduce signaling overhead. For example, in ultra-dense networks, AI can predict the next moment location of vehicles and select BSs according to the predicted location, thereby reducing the switching frequency and switching failure rate and saving the switching signaling overhead [95]. On the other hand, using techniques such as reinforcement learning, AI can dynamically optimize mobility management strategies based on the network environment and user behavior to better cope with complex network environments. Take D2D communication as an example. Traditional cellular networks require BSs to relay user communications, which helps management but adds to transmission latency. The introduction of D2D communication allows direct communication between users within a short distance, improving communication quality and reducing delays. The reuse of cell user resources can improve spectrum utilization [96]. However, this will also interfere with cellular users and increase the management burden of BSs. By utilizing deep reinforcement learning technology, a D2D user terminal functions as an agent capable of autonomously determining the best spectrum access plan without prior knowledge. This can lessen the load on the BS and enhance system efficiency.
3.3 Air Interface Performance Enhancement
Figure 10: Different AI use cases on the air interface. The user will first estimate the downlink channel through the pilot signal, which assists the BS in selecting a good beam to target the user through the feedback channel. Then, communicate with the corresponding user using the selected beam information.
In [97], a white box countermeasure attack evaluation method for the physical layer of semantic communication was proposed for the first time, which generates a high-concealed, data-independent and controllable physical layer attack generation method, solves the problem of 6G semantic communication air interface disturbance generation in a substantial interference wireless environment, and develops a hybrid countermeasure training method. The attack evolution method can counter a variety of typical adversarial interferences and can effectively assist wireless networks in detecting abnormal situations.
The performance of the wireless air interface will directly affect the user’s communication experience. Finding a better air interface algorithm to optimize air interface performance has always been the goal of industry insiders, as shown in Figure 10. However, user demands are increasing with the continuous increase of communication services. Traditional algorithms incur complexity through complex matrix operations and reach the performance bottleneck.
The booming development of AI has provided a new direction for air interface transmission in communication. AI can fit complex and nonlinear mapping relationships. Compared with traditional methods, AI is more suitable for applications in complex time-varying wireless environments. Thus, AI can replace conventional communication methods to improve wireless network performance significantly. In this section, we show how AI technology can improve the transmission performance in OFDM receivers, CSI feedback and beam management.
3.3.1 AI-Based Channel Estimation
Channel estimation [98], [99] is one of the most critical components of wireless communication systems. The wireless environment is complex and time-variant [100][101]; the signals are subjected to a variety of interference in the propagation process; when the signals reach the receiver, the amplitude, phase and frequency of the signals change. The role of channel estimation and channel equalization is to recover the signal as much as possible. In this sense, a good analysis and equalization algorithm is crucial to the receiver’s performance, which determines the final resolution rate of the signal.
Channel estimation algorithms based on training data sequences such as ”least square method” [100],[102] and ”least mean square error” [103],[104], are widely used in channel estimation. The scene of the wireless channel is complex, and the state changes in real time. When estimating the channel in real systems, the parameters in the channel estimation algorithm or the assumed model are prone to deviate from the actual channel states. As a result, the accuracy of the estimated channel can be different from the actual result. AI can remedy the mismatch between mathematical models and real scenarios in this communication [105, 106, 101, 107]. Principle: AI can learn any hidden structure and parameter to fit any complex function. Using AI to perceive wireless scenes and characterize network state space provides a new means of communication. The channel estimation process of actual scenes can be constructed into a neural network, and the algorithm’s performance can be optimized by adjusting the structure and parameters of the neural network.
Considering the sparse mmWave channel matrix as a natural image, the authors of [105] propose a practical and accurate channel estimation framework based on the fast and flexible denoising CNN. Simulation results validate that the training speed of a quick and flexible denoising CNN is more rapid than that of other channel estimators without sacrificing normalized mean square error performance. [106] considered the time-frequency response of a fast-fading communication channel as a 2D image. This scheme considers the pilot values as low-resolution images and uses an SR network cascaded with a denoising IR network to estimate the channel. The results confirm that this pipeline can be used efficiently in channel estimation. To address channel distortion, the model developed in [101] is first trained offline using the data generated from simulation based on channel statistics and then used for recovering the online transmitted data directly. [101] is more robust than conventional methods when fewer training pilots are used, the cyclic prefix is omitted, and nonlinear clipping noise exists. The authors of [107] exploited a learned denoising-based approximate message passing network, which can learn channel structure and estimate channels from many training data. Simulation results show that this neural network significantly outperforms state-of-the-art compressed sensing-based algorithms even when the receiver is equipped with a small number of RF chains.
It can be seen that AI-based channel estimation can combat signal distortion caused by hardware imperfection and wireless channel fading, which narrows the gap between theory and practice. The application of AI to channel estimation has a wide range of application prospects and development potential, and improving the stability and generalization of AI models is also the main research direction in the future.
3.3.2 AI-Based OFDM Receiver
The wireless communication transmitter and receiver are divided into multiple processing mods in OFDM. Each module is responsible for specific sub-tasks, such as Quadrature amplitude modulation, an IFFT, channel estimation, signal detection, etc. The advantage of this design is that each module can be analyzed and optimized individually, resulting in a very efficient and stable system.
However, the performance of the traditional modular OFDM receiver depends on the accuracy of the acquired channel information. Generally speaking, the more specific the channel considered in the algorithm design of each module, the better the receiver’s performance. However, the actual system obtains the channel state information, and the parameters in the channel estimation algorithms or the models are prone to deviate from the actual channel state, resulting in the receiver’s performance being much different from the results obtained.
Channel estimation is one of the most critical technologies in the physical layer. The performance of channel estimation directly affects the quality of communication. Because the performance of the air interface is affected by many environmental factors, the channel estimated by traditional channel estimation methods often has a big gap with the actual channel. AI method can extract essential information in sparse and time-varying pilot frequencies. The complex mapping relationship between the pilot signal and the exact channel is effectively constructed, significantly improving channel estimation accuracy.
Figure 11: This is the difference between traditional OFDM and AI-based OFDM receivers. Using AI modules to replace traditional physical layer processing, AI’s capabilities can compensate for signal distortion.
The mismatch between the mathematical model and the reality can be made up using AI as shown in Figure 11. In theory, AI can learn any hidden structure and parameter to fit any complex function. Using AI to perceive wireless scenes and characterize network state space provides a new means of communication. It can encode the information of actual scenes into neural networks and optimize the performance of algorithms by adjusting the structure and parameters of neural networks.
Regarding an OFDM wireless receiver, some researchers have used a fully connected deep neural network to improve the existing modular OFDM receiver. The process of the OFDM receiver after removing the cyclic prefix is regarded as a black box. Many data samples are obtained as a training set to train the neural network, learn the parameters of each layer of the neural network, and minimize the difference between the output of the neural network and the original information. In this way, the fully connected neural network receiver, after learning, uses a large number of pilot frequency and corresponding channel data as the input of multi-layer fully connected neurons and the corresponding demodulation signal as the output. Following extensive learning, the fully connected neural network receiver employs many pilot frequencies and matching channel data as the input of multi-layer fully connected neurons, with the accompanying demodulation signal serving as the output. Fitting the complex mapping relationship between the two can effectively improve the signal reception performance and demodulation accuracy and has a powerful ability to perform channel estimation and signal detection. Thus, the performance of the receiver is significantly improved.
In [108], a model-driven deep learning (DL) approach was proposed that combines deep learning with expert knowledge to replace existing orthogonal frequency division multiplexing receivers in wireless communications. In [109], a machine learning (ML) assisted physical layer receiver technique was proposed to demodulate the OFDM signals, subject to very high Doppler effects and corresponding distortions in the received signal. To find a balance between full-size CP and non-existent CP, the authors of [110] investigate the redundancy problem and propose a minimum redundant OFDM receiver using deep learning (DL) tools. In [111], the receivers were designed based on deep neural networks (DNNS) in deep learning, consisting of a layer of DNNS and soft decisions. The problems of channel estimation error, delay, and decoding limits between users with traditional detection methods have been solved. It can recover the symbols of all users at one time and jointly perform channel estimation and signal detection. In [112], a novel generative supervised deep neural network is designed that can be trained using a reasonable number of pilots. After channel estimation, the neural network-based receiver (specifically, the autoencoder) jointly learns the pre-encoder and decoder for data symbol detection. In [113], an intelligent receiver for the OFDM communication system was designed based on the Densenet neural network structure, and the intelligent receiver is realized by optimizing the Densenet neural network structure. This method can recover information at the receiving end and avoid complicated pilot operations and signal error accumulation. In [114], a CNN (CNN) was used to reduce the bit error rate by using the mathematical calculation function of discrete Fourier transform (DFT) and the training of OFDM signal samples.
3.3.3 AI-Based Beam Management
Compared with traditional multi-antenna technology, the large-scale multi-antenna system produces more beams and narrow widths and faces new challenges in beam alignment. In September 2022, the IMT 2030 (6G) promotion group released the Research report[115] on Very large MIMO technology, pointing out that hierarchical scanning is generally adopted to align the beam quickly. In the existing academic literature, the beam selection methods of multi-antenna systems can be divided into several categories, including beam selection based on beam scanning[116],[117], beam selection based on training signals[118], beam selection based on position prediction[119],[120], and beam selection[121],[122] based on hierarchical search.
The traditional beam management methods have some limitations when applied to millimeter wave Massive MIMO systems [123]. The primary cause is that the real-world wireless propagation environment often has a variety of it. After a beam is narrowed, the signal transmission is more susceptible to occlusion. The significant increase in the number of beams significantly increases the number of beam search operations in beam alignment. It is difficult to accurately select the optimal beam instantaneously.
AI can fully explore the correlation between the beams themselves, which allows for solving the problems existing in the traditional beam management method as shown in Figure 12. In the actual communication scenario, the massive historical beams provide enough data for training AI models, and the search for the optimal beam based on the correlation between beams constructed by AI dramatically reduces the search space of the beam. At the same time, due to the implicit structure of AI, more complex functional relationships can be explored. Compared with the traditional beam management method received signal strength (RSS), even in a low signal-to-noise ratio environment, the optimal beam selection based on AI still has obvious advantages and robustness. As shown in Fig. 12, red and blue represent antenna arrays with two polarization directions. AI is used to intelligently select beams without beam search, effectively reducing system complexity.
A wide beam can also be used to predict a narrow beam, reducing the beam search space and improving the accuracy of beam prediction. As the input of the AI model, the wide beam is used to predict the narrow beam without activating all antenna arrays, thus reducing the computational search cost on the BS side. Compared with the RSS method, the prediction accuracy is still high in low SNRs.
3.3.4 AI-Based CSI Feedback Algorithm
With the development of multi-antenna technology, CSI feedback design has become a critical point in the research of broadband wireless communication systems. In frequency division duplex mode, the terminal estimates the channel through a downlink pilot to obtain downlink channel information. Since the reciprocity of upstream and downstream channels is not apparent, the terminal needs to feed CSI back to the BS through the uplink. The BS side needs to know the downlink channel to form a beam with better alignment. As the uplink resources are precious, ensuring the accuracy of channel reconstruction and reducing the CSI feedback cost as much as possible is one of the bottleneck problems to overcome in the massive MIMO system based on Frequency Division Duplex.
Traditional CSI feedback includes codebook-based CSI feedback and compressed sensing-based CSI feedback. The standard maker designed the traditional codebook-based CSI feedback in advance and cannot adapt to all possible channel state changes. Sometimes, there is a significant feedback error [124]. CSI feedback based on compressed sensing [125] relies on the sparsity of the channel matrix to ensure performance. In other words, if the channel is not sparse enough, the accuracy of the CSI reconstructed based on compressed sensing theory will be affected. In addition, under the limited feedback overhead, either type of traditional CSI feedback requires many matrix operations, resulting in computing power overhead.
The massive CSI data and the inherent random characteristics of CSI make it preferable to introduce AI to design a new CSI feedback mechanism, as shown in Figure 13 compared with the traditional method. For example, the nonlinear characteristics of deep learning can be used to extract the features of CSI efficiently, and the original CSI data can be transformed into a more compact feature representation, thus reducing the communication overhead in the CSI feedback process. An example of an implementation is the autoencoder architecture, which uses neural networks to extract and compress the features of CSI channel information. The autoencoder consists of an encoder and a decoder. The encoder compiles the high-dimensional channel state information into code words. The decoder is responsible for decoding the code words into the original CSI. The internal structure of encoders and decoders of different algorithms is different, resulting in a difference in channel reconstruction performance. The authors of [126] apply deep learning to CSI feedback first. The proposed CsiNet scheme uses convolution to extract channel features, compress and reconstruct channels, and is superior to the CSI feedback scheme based on compressed sensing regarding feedback accuracy and computational complexity.
The studies in [127, 128, 129, 130] increase the size of the convolutional kernel based on CsiNet to improve the perceptual field of view of the convolutional layer, which is conducive to extracting the sparsity of the channel and further enhancing the accuracy of channel reconstruction. Although a larger perceptive field of view can effectively extract the sparsity of CSI, a smaller convolution kernel can extract finer features from CSI. In [131], a Non-Local neural network is introduced based on the CsiNet network to capture a wide range of dependencies, and the accuracy of channel recovery was improved compared to CsiNet. CRNet proposed in literature [132] uses convolution kernels of different sizes for channel feature extraction and recovery, reducing the computation and improving the accuracy of channel reconstruction.
Figure 13: GCS performance between AI-based CSI feedback and codebook-based CSI feedback.
3.4 New Technological Applications Enabled
AI algorithms can dynamically allocate resources, such as bandwidth and power, to create and manage network slices in the wireless network. This can enable the creation of customized networks tailored to the needs of different applications and services. The booming development of AI has enabled the development of many new technological applications, such as network slicing, mobility management, and wireless localization.
3.4.1 AI-Based Network Slicing
As a core component of 5G technology, network slicing divides infrastructure networks into multiple proprietary logical networks through virtualization technology to meet the customized needs of different vertical industries [133, 134]. Based on isolation principles, customization and end-to-end, network slicing can provide users with differentiated, mutually isolated, and customizable network services with functions and capacities while ensuring the realization of the end-to-end service level agreement (SLA) for services. Although network slicing technology has brought revolutionary changes to 5G networks, its implementation still faces challenges and limitations[135, 136]. Firstly, the complexity of managing and maintaining network slicing increases with the number of slices while meeting the customized needs of different industries and applications. Secondly, the dynamic network environment and changing service demands require network slicing to adapt and adjust quickly, which is often difficult to achieve under the traditional network management framework. In addition, ensuring the isolation between different network slices to prevent resource contention and performance interference is also a vital issue in implementing network slicing [137].
After introducing AI technology, the management and optimization of network slicing will face changes [138][139]. AI technology can automate the configuration, optimization, and troubleshooting of network slicing through a data-driven approach, significantly improving network flexibility and efficiency. For example, AI technology can use raw monitoring data from SDN (software-defined network) controllers to predict the impact of new slicing on the network and make the right resource management decisions to ensure efficient use of network resources while reducing operational and maintenance costs [140]. AI can also enhance the adaptive capability of network slicing, enabling it to respond in real time to changes in network conditions and the evolution of service demand, thus providing more stable and high-quality services.
3.4.2 AI-Based Wireless Localization
Wireless positioning refers to obtaining the location information of mobile users in a cellular environment. This location information is usually in the form of the mobile user’s geographic coordinates relative to some reference points. Wireless location is also commonly referred to as mobile location, wireless location, and geolocation [141].
Wireless positioning technology can be divided into wide and local area positioning according to the scope of services. Wide area positioning provides users with universal positioning services. Local area positioning is mainly used in indoor scenarios. Wide area positioning primarily includes mobile communication network positioning and global satellite navigation systems (e.g., the global positioning system). The GPS is a kind of high-precision radio navigation positioning system based on artificial Earth satellites, which can provide accurate geographical position, vehicle speed and accurate time information anywhere in the world and near-earth space. Local wireless positioning mainly includes Wi-Fi, Bluetooth, UWB, etc. Its positioning technology can be divided into four types: ranging positioning based on signal strength (RSS), fingerprint positioning based on signal characteristics, and positioning based on signal transmission time and Angle measurement [142].
Many applications need high-precision positioning, such as industrial AGV and asset tracking, especially indoor precision positioning, but GPS cannot be used indoors. Despite having a specific positioning reference signal, LTE cannot meet the requirements for high-precision positioning because of its low positioning accuracy and 100-meter BS separation (20 MHz). Bluetooth, Wi-Fi, UWB and other wireless location-based technologies have high deployment costs and are difficult to become a universal positioning technology [143].
AI technology has been introduced to achieve high-precision positioning in the 5G era. One approach is to utilize the random forest technique, which can create a classification model to split a vast area into many small grids and then predict the grid where the user is located. It can also make a regression model to predict the user’s position coordinates. In another method, multi-layer perceptron has a similar application to random forest and can establish regression and classification models to locate targets [144]. AI algorithms can also be trained to recognize the fingerprint of a wireless signal in a specific location and process and fuse data from different sensors (such as Wi-Fi and Bluetooth) to achieve high-precision positioning.
With the widespread use of massive MIMO technology, AI has shown more significant advantages in high-precision positioning [7]. Since systems already use CSI, there is no additional cost for positioning. Using the location in the LOS/NLOS channel co-existence scenario based on the neural network as an example, machine learning and a significant volume of channel data can help effectively map the relationship between channel response and position coordinates, resolving the location problem in complex environments and increasing location accuracy. Very high positioning accuracy is achieved without negligible additional cost, even about 20 mm accuracy under indoor LOS and NLOS conditions.
3.4.3 AI-based Situational Awareness
Network situational awareness technology can bring real-time monitoring, prediction and response capabilities to mobile communication networks, including network management [86] and optimization [87],[84],[89], security monitoring, and fault detection and prediction [145]. The technology includes three levels of environmental element awareness, situation understanding and situation prediction. The network system can recognize and understand various network environmental factors, such as network state, network data, and user behavior patterns while making predictions and timely responses to the future trends of the network state. Figure 14 shows the AI-empowered situational awareness.
Figure 14: AI-empowered situational awareness
In the traditional situational awareness system, the rule engine is typically used to realize the perception of the network state. This sensing method is based on pre-defined rules and logic, which depends on the experience of the rule maker and cannot fully deal with the underlying information and patterns in big data. Traditional systems use data mining and statistical analysis methods to make predictions for network situation prediction models. These methods face problems such as inflexibility and difficulty dealing with complex patterns.
Applying AI technology can significantly enhance the performance of situational awareness systems. Network awareness enabled by AI technology is a data-driven network cognition mode, which can effectively improve the problems existing in traditional situational awareness systems. The massive network data collected by the system under the situational element awareness model can provide data support for AI. In this model, the system can more accurately identify the state of different modes in the network, including traffic, logs, and network KPIs, through AI methods such as deep learning and predict the development trend of the network. At the same time, AI technology can use the perceptual information and the predicted state for analysis and make accurate and effective response decisions.
Specifically, it can be applied from the following four aspects:
1) Application of deep learning in network situation awareness: The application of deep learning in network situation awareness mainly focuses on effectively analyzing and processing large-scale network data to predict network state and future trends accurately [146]. Traditional situational awareness systems rely on pre-defined rules and logic, and these approaches often fail to deal with complex patterns and underlying information in big data. In contrast, by building complex neural network models, deep learning can extract valuable features from massive network traffic and log data, and make efficient classification and prediction. For example, deep learning models like CNN and RNN are used to process network traffic data to identify abnormal behavior and security threats. These models can learn and adjust automatically to adapt to changes in different network environments, thus improving network security monitoring and fault detection efficiency and accuracy.
2) Reinforcement learning application in Ad-hoc network architecture: In the Ad-hoc network architecture, reinforcement learning is widely used to optimize network resource allocation and automate network management[147]. Through trial and error and reward mechanisms, reinforcement learning enables the network to independently adjust and optimize decisions according to the current environment and goals[148]. For example, intelligent agents trained by reinforcement learning algorithms can adjust network configuration parameters in real time under changing network conditions to maximize network performance and user experience. Reinforcement learning algorithms can significantly reduce the need for human intervention, improve the adaptability and fault tolerance of the network, and reduce the cost and complexity of network management.
3) Data-driven network cognition models: A fundamental problem faced by traditional situational awareness systems is that they often rely on human-made rules and static logic, which are difficult to adapt to dynamic and complex network environments. A data-driven network cognition model uses big data technology and machine learning methods to extract knowledge from network data and realize intelligent network management and optimization[149]. By establishing a data-based network cognitive model, the system can collect, analyze and understand various states and changes in the network in real time, such as traffic patterns, network topology and user behavior patterns. In this mode, the system realizes real-time perception of network status, predicts future trends, and automatically adjusts network policies and configurations to improve overall performance and security.
In the future, network situational awareness technology development will face new challenges and opportunities. Key developments include further improving the efficiency and accuracy of deep learning models in network data processing, optimizing the application of reinforcement learning algorithms in complex network environments, and promoting the practice and standardization of data-driven network cognitive models. Multiple technologies and challenges, such as data privacy and security protection, algorithmic adaptability and robustness, must be addressed to achieve these goals.

### v3

@@ 2.2.4 AI-Based Mobility Management
Mobility management is a fundamental function of mobile communication systems, encompassing the network’s management of mobile UE location information, security, and service continuity. Conventional mobility management typically functions by depending on a set of meticulously pre-configured policies and rigid rules. Nevertheless, in practical application contexts, these conventional mobility management strategies have progressively manifested a multitude of bottleneck issues and are incapable of efficiently addressing the intricate and variable 6G network environment as well as the diversified demand circumstances of users. Specifically, when the number of mobile users increases rapidly, these methods may not be able to quickly and effectively handle interactions between users, leading to excessive network load and decreased performance. Second, in complex environments such as large public spaces or major events, conventional mobility management approaches may struggle to cope, resulting in network congestion. Additionally, for high-speed users, such as those in vehicles, frequently hand off can lead to unstable network connections.
To effectively transcend these challenges, AI technology, leveraging its robust data processing capabilities, remarkable adaptive learning mechanisms and highly automated decision-making and optimization capabilities, is integrated into the domain of mobility management. AI is capable of conducting in-depth data mining and analysis of extensive network data, such as user behavior data, network traffic data and channel quality data, to extract valuable insights and latent regular patterns. Additionally, it can automatically and dynamically adjust and continuously optimize mobility management strategies without the need for manual human interference, relying on the acquired knowledge. As early as 2015, 3GPP began researching how AI technology could be integrated into the 5G network. For example, the 3GPP SA2 and SA5 working groups (WG) launched projects on “Enablers for Network Automation” and “Enhancement of Management Data Analytics Services”, which made good progress [89].
Specifically, by means of technologies such as DL, AI is capable of predicting user behaviors including moving speed and direction, thereby permitting the proactive selection and hand off between cells to diminish signal overhead. For example, in ultra-dense network scenarios, frequent cell exchanges can lead to increased overhead and connection disruptions, adversely affecting QoS. The authors of [90] proposed the use of a GRU-based RNN to adaptively handle hand off decisions, significantly reducing hand off latency and improving network throughput. In [91], the authors suggested a method for predicting the next position of vehicles using AI in ultra-dense network scenarios, which allows for BS selection based on predicted positions, thereby reducing hand off frequency and failure rates, and saving hand off signaling overhead. This demonstrates that AI-based mobility management is feasible and practical.
On the other hand, through RL and other techniques, AI can dynamically optimize mobility management strategies based on network environment and user behavior, better addressing complex network environments. As wireless networks at higher frequencies require more dense deployment of BS, the authors of [92] addressed the issue of reduced network capacity due to frequently hand offs, proposing a centralized and multi-agent deep Q-network (DQN) algorithm. Both algorithms can find near-optimal user-to-BS allocations. As shown in Figure 11, the proposed DQN algorithm significantly reduces or even completely avoids ping-pong hand offs compared to the baseline algorithm (conventional SINR-based hand off algorithm). The authors of [93] proposed a ML-based framework to optimize the hand off between sub-6 GHz and millimeter wave bands in vehicular networks, using RL to enable vehicles to predict and discover the optimal hand off target in real time. The proposed method reduces hand off latency, improves connectivity in vehicular networks, achieves seamless hand off even in high mobility scenarios, and enhances user experience.
Figure 11: Ping-pong hand off rate performance comparison in [92]
In light of the aforementioned analysis, We can observe that mobility management is an eternal and significant topic in wireless networks. Most conventional solutions were designed and validated in previous network generations and may not meet future demands for higher communication quality and user experience. AI-based mobility management, including advanced and efficient self-optimization of cell selection, load balancing, and hand off parameters, etc., will play a more critical role in future 6G networks.
2.2.5 AI-Based Resource Orchestration
In the context of the deep fusion between 6G communication networks and the IoT, with hundreds of billions of devices being connected, the generated data is undergoing a continuous explosive growth [94]. Merely strengthening the communication capabilities is insufficient to meet the real-time requirements for data processing in control application scenarios [95]. By harnessing the communication and computing capabilities at the network edge and on the device side, the entire process of data collection, processing, analysis, and decision-making can be accomplished closer to the device side, avoiding the latency drawbacks caused by congestion in the CN, and thus significantly enhancing the data processing response speed. In light of the high-dimensional characteristics and unpredictability of the system resource state space, AI-based resource orchestration strategies (including RL and DRL) can attune to the alterations in the system state via continuous interaction with the system environment, thereby augmenting the overall performance of the system.
The joint scheduling of communication and computing resources founded on RL attains the collaborative and optimal allocation of communication and computing resources via the interaction and learning between the agent and the 6G network environment [96]. Confronted with the resource allocation difficulties within the space-air-ground integrated network in the 6G era, the authors of [97] put forward the RL-assisted bandwidth aware virtual network resource allocation algorithm. This algorithm uses RL with a policy network for node embedding, preferentially handling high-bandwidth requests to meet strict user bandwidth demands. Given the relevant problems that exist within 6G mobile devices and communication systems, motivated by the concept of virtual network embedding, the authors in [98] pioneered a multi-objective aware dynamic resource scheduling algorithm designed for multi-layer computing networks, aiming to augment resource flexibility. They established a scheduling network underpinned by DRL and refined the learning process, thus proffering a sustainable resolution for resource scheduling strategies. Regarding the circumstance that the applicability of the open radio access network (O-RAN) in governing and optimizing the functions of the wireless access network within the 6G network has not been extensively explored, the authors of [99] developed low-complexity algorithms based on methods such as RL to address the problems of jointly optimizing traffic splitting and allocation, congestion control, and scheduling across different time scales. This effort provides insights for realizing a fully automated network with enhanced control and flexibility.
On the other hand, the dynamic scheduling of network slices based on DRL exploits the potent feature extraction and learning capabilities of DNNs to manage network slice resources intelligently. To address the challenges of computing resource demands and intelligent resource allocation in 6G network IoT services, the authors of [100] proposed an effective DRL-based solution for IoT resource expansion and service placement, and verified the effectiveness of its multi-application autonomous resource allocation via dataset simulation. According to [101], the authors proposed a multi-task DRL method based on graph convolutional networks by introducing a joint network slicing and routing mechanism, achieving robustness in various network environments. In the study of [102], the authors established a model-free DRL framework by a hierarchical structure integrating modified deep deterministic policy gradient (DDPG) and double DQN to actualize an intelligent RAN slicing strategy with two-layer control granularity for maximizing QoS and slice spectrum efficiency. Moreover, a dynamic spectrum allocation scheme was proposed in [103], which utilized the advantages of the DQN and reduced the search state explosion by a reward and punishment framework to dynamically allocate unallocated resource blocks to mobile units and obtained the Pareto optimal solution of sub-problems via Chebyshev decomposition.
The AI-based resource orchestration has already demonstrated certain advantages in fields such as 6G and has achieved remarkable results in resource allocation and utilization by combining various techniques. However, it still faces numerous challenges and opportunities in the future. On the one hand, efforts should be focused on optimizing and innovating algorithms. For instance, the improvement of DRL algorithms is urgent in order to cope with complex and changeable environments. On the other hand, when deepening the integration of multiple technologies, not only should the combination paths with emerging technologies be actively explored, but also the compatibility issues during the integration process should be addressed. Meanwhile, it is necessary to strive to enhance the adaptability to complex scenarios, meet the diverse resource requirements in various special scenarios of different industries, and improve performance evaluation and enhance algorithm interpretability, thereby achieving more extensive applications and development.
2.2.6 AI-based Situational Awareness and Fault Detection
Network situational awareness technology is capable of endowing mobile communication networks with real-time monitoring, prediction, and response capabilities. Its application range spans various aspects, including network management and optimization, security monitoring, as well as fault detection and prediction [104]. The AI-based situational awareness system can ensure the security of all links within the 6G network. It can prevent malicious penetration attacks, data loss, and malicious tampering, guarantee the credibility of access terminals, the reliability of transmission channels, and the controllability of service applications, and achieve panoramic security monitoring, thereby comprehensively enhancing the comprehensive security defense capabilities of all 6G services.
Specifically, the situational awareness technology relying on AI is a data-driven network cognitive paradigm. It is capable of collecting, storing, and analyzing vast amounts of network data via big data platforms and DL techniques, thereby providing precise threat intelligence and full traffic inspection [105]. This technology employs AI approaches such as DL to precisely identify the states of various elements including network traffic, logs, and network KPI, predict network development trends, and formulate accurate response strategies through the analysis of perceived information and predicted states [106], realizing the shift from “passive defense” to “active defense”. The AI-driven threat representation and causal reasoning can automatically identify and infer the origins of attacks, preventing the recurrence of similar security events. The real-time threat alert and response handling capabilities assist users in quickly grasping attack circumstances, affected devices, and emergency handling tactics, thus ensuring the visualization and continuous monitoring of network security postures.
The network fault analysis constitutes one of the crucial elements in the O&M of mobile communication networks. As the number of mobile communication access devices increases and the network scale expands, the quantity of network alarms is growing explosively, and the relationships among them are highly intricate [107]. Employing AI to build a fault detection analysis model endows the model with the capacity for high-dimensional data analysis, enabling it to serve as a more intelligent and efficient tool for fault analysis [108]. The fault detection method based on AI automatically extracts key features from massive network alarm data through DL and ML algorithms. It combines network topology structures, configuration settings, KPI, and historical alarm and fault handling data to construct a fault analysis model. Through conducting a fusion analysis of these data, the system can effectively identify common issues within the network and perform accurate fault diagnoses. To enhance the system’s adaptability and precision, the results obtained after fault handling will be utilized as feedback to update and optimize the fault diagnosis model, allowing the system to operate efficiently under constantly changing network environments and business requirements [109]. The specific operational workflow is illustrated in Figure 12.
Figure 12: AI-based fault detection
By integrating AI technology, we can establish a systematic and intelligent network fault analysis and tracing back scheme, which can provide automated, accurate and flexible fault detection and tracing back capabilities and continuously enhance the efficiency of fault diagnosis and handling. Meanwhile, it can also adapt to the changes of network environment and business requirements.